OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of x2 = 0. I need some help. A. no real roots B. one real root C. two distinct real roots D. three distinct real roots

OpenStudy (anonymous):

is it x^2=0 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

this has two equal roots.....0,0

OpenStudy (anonymous):

so the answer is c?

hartnn (hartnn):

do you know what discriminant is ?

OpenStudy (anonymous):

no that's why im asking what the answer is so i can figure out why it's the answer

hartnn (hartnn):

Compare your quadratic equation \(x^2=0\)with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] Then your discriminant is \(D = b^2-4ac \)

OpenStudy (anonymous):

that's where it gets confusing idk how to figure out what the root would be

hartnn (hartnn):

First step is to find, a,b,c values. can you do that ? or need help with that ?

OpenStudy (anonymous):

lol i don't even know how to do that. i suck at math

hartnn (hartnn):

 now can you tell me values of a,b,c ?